if Tan8A - Tan5A - Tan3A = k . Tan8A .Tan5A.Tan3A find k

Grade:11

1 Answers

Harshit Singh
2 years ago
Dear Student

tan8A –( tan5A + tan3A) =k . Tan8A .Tan5A.Tan3A
tan(3A + 5A ) = tan8A = tan3A + tan5A/(1-tan3Atan5A)
tan5A + tan3A= tan8A * ( 1 –tan3Atan5A)
tan8A –( tan5A + tan3A)= tan8A –(tan8A * ( 1 –tan3Atan5A))
= tan8A –tan8A + tan8Atan3Atan5A
= tan8Atan3Atan5A
hence , k=1

Thanks